OpenStudy (anonymous):

a total of 411 tickets were sold for the school play. They were either adult tickets or student tickets. The number of student tickets sold was two times the number of afult tickets sold. How many adult tickets were sold?

OpenStudy (anonymous):

102.5? adult tickets?

OpenStudy (anonymous):

x=adult tickets hence student tickets = 2x 2x+x=411, 3x=411, x= 137, student tickets = 2*137

hero (hero):

s = student a = adults s + a = 411 s = 2a Solve the system

OpenStudy (anonymous):

Well, if the number of student tickets was two times the # of adult tickets, that means that, since total number of tickets = 411=# child tickets+# adult tickets, 411 = 2*# of adult tickets + # of adult tickets (I just substituted 2*# of adult tickets for # child tickets) This means that 411 = 3*# of adult tickets. If you divide both sides by 3, you get # of adult tickets = 137

OpenStudy (anonymous):

Thank you!